# Padé Approximant Neural Networks for Enhanced Electric Motor Fault Diagnosis Using Vibration and Acoustic Data

### v1

@@ Abstract
Purpose The primary aim of this study is to enhance fault diagnosis in induction machines by leveraging the Padé Approximant Neuron (PAON) model. While accelerometers and microphones are standard in motor condition monitoring, deep learning models with nonlinear neuron architectures offer promising improvements in diagnostic performance. This research addresses the question: Can Padé Approximant Neural Networks (PadéNets) outperform conventional Convolutional Neural Networks (CNNs) and Self-Organized Operational Neural Networks (Self-ONNs) in diagnosing electrical and mechanical faults using vibration and acoustic data?
@@ 1 Introduction
While CNNs exhibit strong performance under controlled conditions, prior studies [24, 25] emphasize that conventional CNNs, built upon a fixed architecture and first-order neuron model, often struggle to capture highly nonlinear and complex patterns inherent in real-world data. Although nonlinearity is introduced through pointwise activation functions, such as ReLU [32] and its variant leaky ReLU [30], these functions are predefined and uniformly applied across layers, limiting the network’s representational flexibility. To address this constraint, Padé Activation Units (PAUs) were introduced in [31] as a learnable alternative to traditional hand-crafted activations. PAUs model activation functions using Padé approximants, which are rational functions expressed as the ratio of two polynomials. This formulation enables the network to learn complex, task-specific nonlinear mappings during training, as both the numerator and denominator coefficients are optimized via backpropagation (BP). By making the activation functions adaptive rather than fixed, PAUs provide a more expressive and flexible framework for capturing intricate data patterns. To further extend this paradigm and enhance the network’s ability to model nonlinearities at the neuron level, Self-Organized Operational Neural Networks (Self-ONNs) have been proposed [27]. Unlike traditional CNNs, which rely solely on pointwise nonlinear activations, Self-ONNs incorporate nonlinear neuron models. They incorporate generative neurons that approximate the necessary nonlinear mappings by utilizing a truncated Taylor series expansion centered at the origin, specifically applying a Maclaurin series expansion up to a predefined order. Therefore, generative neurons use the input along with its higher-order powers and compute their weighted sum to approximate a nonlinear mapping in the neuron itself. The enhanced fault diagnosis performance of 1D and 2D Self-ONNs has been validated in studies on machinery fault diagnosis, utilizing various sensor modalities [14, 15, 20, 22]. While generative neurons in Self-ONNs capture greater nonlinearity, the linear combination of different input orders can lead to instability outside a safe computation range. Moreover, since Taylor series approximations are most accurate near the expansion point, the output of generative neurons is constrained by a tanh activation function, which may suffer from vanishing gradients during training. To address this limitation, a new class of networks known as PadéNets, built using Padé Approximant Neurons (PAONs) has been proposed [18]. Padé neurons utilize Padé approximation at the neuron level by representing nonlinear functions as ratios of polynomials. In PadéNets, a single neuron within a 1D Padé layer with a kernel size of $k$ effectively learns $k$ distinct Padé approximants, each represented as a ratio of two polynomials. This structure significantly increases the degrees of freedom available to the model compared to PAUs, as it introduces nonlinearity within the kernel itself, in addition to the nonlinearity contributed by the activation function. They have been shown to offer improved performance compared to Taylor-based generative neurons and convolutional neurons in single-image super-resolution tasks [18]. Furthermore, PAONs generalize several existing neuron models and can effectively serve as a replacement for conventional convolutional neurons within CNN architectures. To leverage the superior feature extraction capabilities of PAONs, this study introduces the use of 1D PadéNets for the classification of electrical and mechanical faults in three-phase induction machines. The main contributions of this work can be summarized as follows:
@@
[…]
@@ 2 Methods
$x_{k}^{l}=b_{k}^{l}+\sum_{i=0}^{N_{l-1}}x_{ik}^{l}$ (1)
In this expression, $x_{ik}^{l}\in\mathbb{R}^{M}$ denotes the feature map obtained by convolving the $i^{\text{th}}$ output map from layer $(l-1)$, denoted as $y_{i}^{(l-1)}\in\mathbb{R}^{M}$, with the kernel $w_{ik}^{l}\in\mathbb{R}^{K}$, which connects it to the $k^{\text{th}}$ input feature map in layer $l$. The term $b_{k}^{l}$ represents the bias associated with the $k^{\text{th}}$ neuron in the current layer. The 1D convolution operation used to compute $x_{ik}^{l}(m)$ is given by:
$x_{ik}^{l}(m)=\sum_{r=0}^{K-1}w_{ik}^{l}(r)y_{i}^{(l-1)}(m+r)$ (2)
In forward pass, each input feature map $x_{k}^{l}$ is then transformed by a nonlinear activation function, resulting in the output feature representation of the convolutional neuron.
@@
$\overline{x}_{ik}^{l}(m)=P_{k}^{l}\left(\psi_{k}^{l}\left(w_{ik}^{l}(r),\;y_{i}^{(l-1)}(m+r)\right)_{r=0}^{K-1}\right),$ (3)
where $\psi_{l}^{k}(\cdot):\mathbb{R}^{M\times K}\rightarrow\mathbb{R}^{M\times K}$ and $P_{k}^{l}(\cdot):\mathbb{R}^{K}\rightarrow\mathbb{R}^{1}$ are called nodal and pool operators, respectively, assigned to the $k^{\text{th}}$ neuron of the $l^{\text{th}}$ layer.
@@
$\tilde{\psi}_{k}^{l}\left(w_{ik}^{l(P)}(r),y_{i}^{(l-1)}(m+r)\right)$ (4)
$=\sum_{p=1}^{P}w_{ik}^{l(P)}(r,p)\left(y_{i}^{(l-1)}(m+r)\right)^{p}$
In Equation 4, the hyperparameter $P$ sets the order of the Taylor polynomial approximation, thereby influencing the degree of nonlinearity. Additionally, the weights $w_{ik}^{l(P)}$ now consist of $P$ times the number of learnable parameters in the convolutional model. During training, the weights $w_{ik}^{l(P)}$ are updated via the standard BP algorithm, leading to nonlinear transformations [27].
@@
$\tilde{x}_{ik}^{l}=\sum_{p=1}^{P}\operatorname{Conv}1D\left(w_{ik}^{l(P)},\left(y_{i}^{(l-1)}\right)^{p}\right)$ (5)
Thus, the formulation can be implemented using $P$ 1D convolution operations. When $P=1$, it simplifies to the conventional 1D convolution. Self-ONNs are super set of CNNs ($P=1$), and generative neurons in Self-ONNs enable the modeling of more complex nonlinearities. However, using higher-order powers of the input can introduce numerical instability outside a stable computational range; therefore, neuron outputs are typically constrained by bounded activation functions. Since Taylor series approximations are most accurate near the expansion point, the output of generative neurons is typically constrained by the tanh activation function to model the nonlinear mapping around the origin. However, the tanh function saturates, which can lead to the vanishing gradient problem during BP, hindering effective training. To address these challenges, a new neuron model, Padé Approximant Neurons (PAONs), inspired by Padé approximants, has recently been proposed [18].
@@
If we let $f_{[P/Q]}(x)$ denote the Padé approximation of a function $f(x)$, where the numerator is a polynomial of degree $P$ and the denominator is a polynomial of degree $Q$, the approximation can be formulated as follows:
$f_{[P/Q]}(x)$ $=\frac{R_{P}(x)}{S_{Q}(x)}=\frac{\displaystyle\sum_{m=0}^{P}a_{m}x^{m}}{\displaystyle\sum_{n=0}^{Q}b_{n}x^{n}}$ (6)
$=\frac{a_{0}+a_{1}x+\cdots+a_{P}x^{P}}{b_{0}+b_{1}x+\cdots+b_{Q}x^{Q}}$
@@
$f_{[P/Q]}(x)=\dfrac{\displaystyle\sum_{m=0}^{P}a_{m}x^{m}}{1+\displaystyle\sum_{n=1}^{Q}b_{n}x^{n}}=\dfrac{a_{0}+\displaystyle\sum_{m=1}^{P}a_{m}x^{m}}{1+\displaystyle\sum_{n=1}^{Q}b_{n}x^{n}}$ (7)
If we interpret the coefficients $a_{m}$ and $b_{n}$ as kernels in a convolution operation, with $a_{0}$ representing the bias, this formulation corresponds to the following Padé neuron model:
$f_{[P/Q]}(x)=\dfrac{w_{p0}+\displaystyle\sum_{m=1}^{P}w_{pm}\ast x^{m}}{1+\displaystyle\sum_{n=1}^{Q}w_{qn}\ast x^{n}},$ (8)
@@
where $w_{pm}$ and $w_{qn}$ represent the 1D convolutional filters for the numerator and denominator, respectively, associated with the input orders $m$ and $n$. Additionally, $w_{p0}$ serves as the bias term in the numerator. One important consideration with this neuron model is that the denominator can potentially become zero or approach to zero throughout training. To mathematically ensure that the denominator remains nonzero, several variants of the Padé neurons have been proposed [18]. In this study, we adopt the first variant, which involves taking the absolute value of each term in the denominator to guarantee that each element in the numerator is divided by a value greater than or equal to one. Therefore, we can express the equation for this variant as follows:
$f_{[P/Q]}(x)=\dfrac{w_{p0}+\displaystyle\sum_{m=1}^{P}w_{pm}\ast x^{m}}{1+\displaystyle\sum_{n=1}^{Q}\left|w_{qn}\ast x^{n}\right|}$ (9)
@@
Each kernel element in a Padé neuron adapts independently, allowing each weight group to learn its own specific Padé approximation. This self-adjustment enhances the model’s nonlinearity by integrating higher-order features in both the numerator and denominator of the approximation. Additionally, as the Padé neuron is expressed as a ratio of polynomials, it provides greater stability, even with higher-order approximations. Thus, there is no need to restrict its output using bounded activation functions like tanh.
@@
Padé neurons generalize both convolutional and generative neuron models. For $P=1$ and $Q=0$, the Padé neuron reduces to a standard convolutional neuron, and for $P\geq 2$ and $Q=0$, they behave as generative neurons. As a result, PAONs can effectively capture complex nonlinear relationships and they are capable of replacing existing neuron models in CNNs and Self-ONNs. Figure 2 offers a comparative illustration of the computations involved in a convolutional, generative, and Padé neuron.

### v2

@@ Abstract
Purpose The primary aim of this study is to enhance fault diagnosis in induction machines by leveraging the Padé Approximant Neuron (PAON) model. While accelerometers and microphones are standard in motor condition monitoring, deep learning models with nonlinear neuron architectures offer promising improvements in diagnostic performance. This research investigates whether Padé Approximant Neural Networks (PadéNets) can outperform conventional Convolutional Neural Networks (CNNs) and Self-Organized Operational Neural Networks (Self-ONNs) in the diagnosis of electrical and mechanical faults from vibration and acoustic data.
@@ 1 Introduction
While CNNs exhibit strong performance under controlled conditions, prior studies [24, 25] emphasize that conventional CNNs, built upon a fixed architecture and first-order neuron model, often struggle to capture highly nonlinear and complex patterns inherent in real-world data. Although nonlinearity is introduced through pointwise activation functions, such as ReLU [32] and its variant $\mathrm{LeakyReLU}$ [30], these functions are predefined and uniformly applied across layers, limiting the network’s representational flexibility. To address this constraint, Padé Activation Units (PAUs) were introduced in [31] as a learnable alternative to traditional hand-crafted activations. PAUs model activation functions using Padé approximants, which are rational functions expressed as the ratio of two polynomials. This formulation enables the network to learn complex, task-specific nonlinear mappings during training, as both the numerator and denominator coefficients are optimized via backpropagation (BP). By making the activation functions adaptive rather than fixed, PAUs provide a more expressive and flexible framework for capturing intricate data patterns. To further extend this paradigm and enhance the network’s ability to model nonlinearities at the neuron level, Self-Organized Operational Neural Networks (Self-ONNs) have been proposed [27]. Unlike traditional CNNs, which rely solely on pointwise nonlinear activations, Self-ONNs incorporate nonlinear neuron models. They incorporate generative neurons that approximate the necessary nonlinear mappings by utilizing a truncated Taylor series expansion centered at the origin, specifically applying a Maclaurin series expansion up to a predefined order. Therefore, generative neurons use the input along with its higher-order powers and compute their weighted sum to approximate a nonlinear mapping in the neuron itself. The enhanced fault diagnosis performance of 1D and 2D Self-ONNs has been validated in studies on machinery fault diagnosis, utilizing various sensor modalities [14, 15, 20, 22]. While generative neurons in Self-ONNs capture greater nonlinearity, the linear combination of different input orders can lead to instability outside a safe computation range. Moreover, since Taylor series approximations are most accurate near the expansion point, the output of generative neurons is constrained by a $\tanh$ activation function, which may suffer from vanishing gradients during training. To address this limitation, a new class of networks known as PadéNets, built using Padé Approximant Neurons (PAONs) has been proposed [18]. Padé neurons utilize Padé approximation at the neuron level by representing nonlinear functions as ratios of polynomials. In PadéNets, a single neuron within a 1D Padé layer with a kernel size of $k$ effectively learns $k$ distinct Padé approximants, each represented as a ratio of two polynomials. This structure significantly increases the degrees of freedom available to the model compared to PAUs, as it introduces nonlinearity within the kernel itself, in addition to the nonlinearity contributed by the activation function. They have been shown to offer improved performance compared to Taylor-based generative neurons and convolutional neurons in single-image super-resolution tasks [18]. Furthermore, PAONs generalize several existing neuron models and can effectively serve as a replacement for conventional convolutional neurons within CNN architectures. To leverage the superior feature extraction capabilities of PAONs, this study introduces the use of 1D PadéNets for the classification of electrical and mechanical faults in three-phase induction machines. The main contributions of this work can be summarized as follows:
@@
Figure 1: The proposed 1D PadéNet-based framework and the diagram of a Padé neuron with $P=2,Q=1$, where $w_{0}$ represents the bias term in the numerator, $(\cdot)^{n}$ denotes element-wise exponentiation of the input to the $n^{\text{th}}$ power, $\mathbf{\ast}$ indicates convolution, and $\frac{\displaystyle\cdot}{\displaystyle\cdot}$ implements Equation 10.
@@ 2 Methods
$\mathbf{x}_{k}^{(l)}=b_{k}^{(l)}+\sum_{i=1}^{N_{l-1}}\mathbf{x}_{ik}^{(l)}$ (1)
In this expression, $\mathbf{x}_{ik}^{(l)}\in\mathbb{R}^{M}$ denotes the feature map obtained by convolving the $i^{\text{th}}$ output map from layer (l1)(l-1), denoted as $\mathbf{y}_{i}^{(l-1)}\in\mathbb{R}^{M}$, with the kernel $\mathbf{w}_{ik}^{(l)}\in\mathbb{R}^{K}$, which connects it to the $k^{\text{th}}$ input feature map in layer $l$. The term $b_{k}^{(l)}$ represents the bias associated with the $k^{\text{th}}$ neuron in the current layer, and $N_{l-1}$ is the number of output feature maps (or channels) produced by layer l1l-1. The 1D convolution operation used to compute $\mathbf{x}_{ik}^{(l)}[m]$ is given by:
$\mathbf{x}_{ik}^{(l)}[m]=\sum_{r=0}^{K-1}w_{ik}^{(l)}[r]\,y_{i}^{(l-1)}[m+r]$ (2)
In the forward pass, each input feature map $\mathbf{x}_{k}^{(l)}$ is then transformed by a nonlinear activation function followed by an optional subsampling operation, resulting in the output feature representation of the convolutional neuron.
@@
$\overline{\mathbf{x}}_{ik}^{(l)}[m]=P_{k}^{(l)}\!\left(\psi_{k}^{(l)}\!\left(w_{ik}^{(l)}[r],\;y_{i}^{(l-1)}[m+r]\right)_{r=0}^{K-1}\right)$ (3)
where $\boldsymbol{\psi}_{k}^{(l)}(\cdot):\mathbb{R}^{M\times K}\rightarrow\mathbb{R}^{M\times K}$ and $P_{k}^{(l)}(\cdot):\mathbb{R}^{K}\rightarrow\mathbb{R}^{1}$ are called nodal and pool operators, respectively, assigned to the $k^{\text{th}}$ neuron of the $l^{\text{th}}$ layer.
@@
$\tilde{\boldsymbol{\psi}}_{k}^{(l)}\!\left(\{\mathbf{w}_{p,ik}^{(l)}[r]\}_{p=1}^{P},\,\mathbf{y}_{i}^{(l-1)}[m+r]\right)$ (4)
$\quad=\sum_{p=1}^{P}\mathbf{w}_{p,ik}^{(l)}[r]\;\big(\mathbf{y}_{i}^{(l-1)}[m+r]\big)^{p}$
In Equation 4, the hyperparameter $P$ sets the order of the Taylor polynomial approximation, thereby influencing the degree of nonlinearity. Additionally, the weights $\mathbf{w}_{p,ik}^{(l)}$ now consist of $P$ times the number of learnable parameters in the corresponding convolutional model. During training, the weights $\mathbf{w}_{p,ik}^{(l)}$ are updated via the standard BP algorithm, leading to nonlinear transformations [27].
@@
$\tilde{\mathbf{x}}_{ik}^{(l)}=\sum_{p=1}^{P}\operatorname{Conv1D}\!\left(\mathbf{w}_{p,ik}^{(l)},\;\big(\mathbf{y}_{i}^{(l-1)}\big)^{p}\right)$ (5)
Thus, the formulation can be implemented using $P$ 1D convolution operations. When $P=1$, it simplifies to the conventional 1D convolution. Self-ONNs are super set of CNNs ($P=1$), and generative neurons in Self-ONNs enable the modeling of more complex nonlinearities. However, using higher-order powers $\big(\mathbf{y}_{i}^{(l-1)}\big)^{p}$ can introduce numerical instability outside a stable computational range; therefore, $\tilde{\mathbf{x}}_{ik}^{(l)}$ is typically constrained by bounded activation functions. Since Taylor series approximations are most accurate near the expansion point, the output of generative neurons is typically constrained by the $\tanh$ activation function to model the nonlinear mapping around the origin. However, the $\tanh$ function saturates, which can lead to the vanishing gradient problem during BP, hindering effective training. To address these challenges, a new neuron model, Padé Approximant Neurons (PAONs), inspired by Padé approximants, has recently been proposed [18].
@@
If we let $f_{[P/Q]}(y)$ denote the Padé approximation of a function $f(y)$, where the numerator is a polynomial of degree $P$ and the denominator is a polynomial of degree $Q$, the approximation can be formulated as follows:
$f_{[P/Q]}(y)=\frac{R_{P}(y)}{S_{Q}(y)}=\frac{\displaystyle\sum_{m=0}^{P}a_{m}\,y^{m}}{\displaystyle\sum_{n=0}^{Q}b_{n}\,y^{n}}$ (6)
@@
$f_{[P/Q]}(y)=\frac{\displaystyle a_{0}+\sum_{m=1}^{P}a_{m}\,y^{m}}{\displaystyle 1+\sum_{n=1}^{Q}b_{n}\,y^{n}}$ (7)
Padé Activation Units (PAUs) adapt the classical Padé approximation for use as learnable activation functions in neural networks [31]. Unlike fixed-form nonlinearities such as $\tanh$, a PAU represents the activation as the ratio of two polynomials whose coefficients are trainable parameters. By appropriately learning these coefficients, PAUs can replicate common activation functions (e.g., sigmoid, $\tanh$) as special cases or generate entirely new, data-driven nonlinearities. The analytical differentiability of the rational form ensures compatibility with standard BP, while its flexibility enables the network to capture complex, task-specific behaviors that may be inaccessible to conventional fixed activations. Formally, a PAU in the $l^{\text{th}}$ layer, applied element-wise to each input map $\mathbf{x}_{i}^{(l)}\in\mathbb{R}^{M}$, is defined as:
$\boldsymbol{\phi}_{\text{PAU}}\!\left(\mathbf{x}_{i}^{(l)}\right)=\frac{\displaystyle\sum_{m=0}^{P}a_{m}^{(l)}\,\big(\mathbf{x}_{i}^{(l)}\big)^{m}}{\displaystyle 1+\sum_{n=1}^{Q}b_{n}^{(l)}\,\big(\mathbf{x}_{i}^{(l)}\big)^{n}},$ (8)
where $a_{m}^{(l)}$ and $b_{n}^{(l)}$ are trainable scalar coefficients shared by all feature maps in the $l^{\text{th}}$ layer.
This idea of embedding trainable Padé approximants into the network structure naturally motivates the use of PAONs in PadéNets, where the polynomial ratio formulation is incorporated directly into the neuron instead of the activation function. If we interpret the coefficients $a_{m}$ and $b_{n}$ as kernels in a convolution operation, with $a_{0}$ representing the bias, the $k^{\text{th}}$ input feature map in the $l^{\text{th}}$ Padé layer can be expressed as:
$\mathbf{x}_{k}^{(l)}=\frac{w_{p0,k}^{(l)}+\displaystyle\sum_{m=1}^{P}\sum_{i=1}^{N_{l-1}}\mathbf{w}_{pm,ik}^{(l)}\ast\big(\mathbf{y}_{i}^{(l-1)}\big)^{m}}{1+\displaystyle\sum_{n=1}^{Q}\sum_{i=1}^{N_{l-1}}\mathbf{w}_{qn,ik}^{(l)}\ast\big(\mathbf{y}_{i}^{(l-1)}\big)^{n}}$ (9)
@@
where $\mathbf{y}_{i}^{(l-1)}\in\mathbb{R}^{M}$ is the $i^{\text{th}}$ output feature map from the $(l-1)^{\text{th}}$ layer, $\mathbf{w}_{pm,ik}^{(l)},\,\mathbf{w}_{qn,ik}^{(l)}\in\mathbb{R}^{K}$ are the numerator and denominator kernels corresponding to polynomial orders $m$ and $n$, respectively, with $i$ indexing the output feature map from layer l1l-1 and $k$ indexing the input feature map in layer $l$. $w_{p0,k}^{(l)}$ is the bias term, and $\ast$ denotes 1D convolution. One important consideration with this neuron model is that the denominator can potentially become zero or approach to zero throughout training. To mathematically ensure that the denominator remains nonzero, several variants of the Padé neurons have been proposed [18]. In this study, we adopt the first variant, which involves taking the absolute value of each term in the denominator to guarantee that each element in the numerator is divided by a value greater than or equal to one. Therefore, we can express the equation for this variant as follows:
$\mathbf{x}_{k}^{(l)}=\frac{w_{p0,k}^{(l)}+\displaystyle\sum_{m=1}^{P}\sum_{i=1}^{N_{l-1}}\mathbf{w}_{pm,ik}^{(l)}\ast\big(\mathbf{y}_{i}^{(l-1)}\big)^{m}}{1+\displaystyle\sum_{n=1}^{Q}\sum_{i=1}^{N_{l-1}}\big|\,\mathbf{w}_{qn,ik}^{(l)}\ast\big(\mathbf{y}_{i}^{(l-1)}\big)^{n}\,\big|}$ (10)
@@
Each kernel element in a Padé neuron adapts independently, allowing each weight group to learn its own specific Padé approximation. This self-adjustment enhances the model’s nonlinearity by integrating higher-order features in both the numerator and denominator of the approximation. Additionally, as the Padé neuron is expressed as a ratio of polynomials, it provides greater stability, even with higher-order approximations. When the numerator and denominator degrees are comparable, the PAON’s rational form maintains training stability even with unbounded activation functions.
@@
Padé neurons generalize both convolutional and generative neuron models. For $P=1$ and $Q=0$, the Padé neuron reduces to a standard convolutional neuron in CNNs, and for $P\geq 2$ and $Q=0$, they behave as generative neurons in Self-ONNs. As a result, PAONs can effectively capture complex nonlinear relationships and they are capable of replacing existing neuron models in CNNs and Self-ONNs. Compared to a standard Conv1D layer, the PAON formulation introduces $(P+Q-1)\times K\times C_{\text{in}}\times C_{\text{out}}$ additional trainable parameters for a kernel size $K$, where $C_{\text{in}}$ and $C_{\text{out}}$ denote the number of input and output feature maps, respectively. This increase is due to the full PAON mapping that consists of $(P+Q)$ parallel convolutional branches, whereas a conventional convolution employs only a single branch. Figure 2 offers a comparative illustration of the computations involved in a convolutional, generative, and Padé neuron.
To enable end-to-end training of networks containing Padé neurons (PAONs), we derive the gradients of the loss function $\mathcal{L}$ with respect to the numerator kernels $\mathbf{w}_{pm,ik}^{(l)}$, the denominator kernels $\mathbf{w}_{qn,ik}^{(l)}$, and the output feature map of the previous layer $\mathbf{y}_{i}^{(l-1)}$. We denote the upstream gradient from the $(l+1)^{\mathrm{th}}$ layer as follows;
$\boldsymbol{\delta}_{k}^{(l)}\triangleq\frac{\partial\mathcal{L}}{\partial\mathbf{x}_{k}^{(l)}},$ (11)
For compactness, we define:
$\mathbf{R}_{k}^{(l)}$ $=w_{p0,k}^{(l)}+\sum_{m=1}^{P}\sum_{i=1}^{N_{l-1}}\mathbf{w}_{pm,ik}^{(l)}\ast\big(\mathbf{y}_{i}^{(l-1)}\big)^{m},$ (12)
$\mathbf{S}_{k}^{(l)}$ $=1+\sum_{n=1}^{Q}\sum_{i=1}^{N_{l-1}}\big|\,\mathbf{w}_{qn,ik}^{(l)}\ast\big(\mathbf{y}_{i}^{(l-1)}\big)^{n}\,\big|,$ (13)
[…]
Since $\mathbf{w}_{pm,ik}^{(l)}$ appears only in $\mathbf{R}_{k}^{(l)}$, differentiating $\mathbf{x}_{k}^{(l)}=\mathbf{R}_{k}^{(l)}/\mathbf{S}_{k}^{(l)}$ with respect to $\mathbf{w}_{pm,ik}^{(l)}$ while treating $\mathbf{S}_{k}^{(l)}$ as constant yields;
$\frac{\partial\mathcal{L}}{\partial\mathbf{w}_{pm,ik}^{(l)}}=\left(\boldsymbol{\delta}_{k}^{(l)}\odot[\mathbf{S}_{k}^{(l)}]^{-1}\right)\ast_{\mathrm{grad}}\big(\mathbf{y}_{i}^{(l-1)}\big)^{m}.$ (14)
For denominator kernels, we introduce $\mathbf{h}_{qn,ik}^{(l)}\triangleq\mathbf{w}_{qn,ik}^{(l)}\ast\big(\mathbf{y}_{i}^{(l-1)}\big)^{n}$ and use the subgradient $\operatorname{sgn}(\cdot)$ element-wise. Applying the quotient rule gives:
$\frac{\partial\mathcal{L}}{\partial\mathbf{w}_{qn,ik}^{(l)}}$ $=\left(\boldsymbol{\delta}_{k}^{(l)}\odot\big(-\,\mathbf{R}_{k}^{(l)}\oslash[\mathbf{S}_{k}^{(l)}]^{\circ 2}\big)\right)$
$\quad\ast_{\mathrm{grad}}\left(\operatorname{sgn}(\mathbf{h}_{qn,ik}^{(l)})\odot\big(\mathbf{y}_{i}^{(l-1)}\big)^{n}\right).$ (15)
For the input gradients, both $\mathbf{R}_{k}^{(l)}$ and $\mathbf{S}_{k}^{(l)}$ depend on $\mathbf{y}_{i}^{(l-1)}$, leading to:
$\frac{\partial\mathbf{R}_{k}^{(l)}}{\partial\mathbf{y}_{i}^{(l-1)}}$ $=\sum_{m=1}^{P}\tilde{\mathbf{w}}_{pm,ik}^{(l)}\ast\big(m\,(\mathbf{y}_{i}^{(l-1)})^{m-1}\big),$ (16)
$\frac{\partial\mathbf{S}_{k}^{(l)}}{\partial\mathbf{y}_{i}^{(l-1)}}$ $=\sum_{n=1}^{Q}\operatorname{sgn}\!\big(\mathbf{h}_{qn,ik}^{(l)}\big)\odot$
$\quad\left[\,\tilde{\mathbf{w}}_{qn,ik}^{(l)}\ast\big(n\,(\mathbf{y}_{i}^{(l-1)})^{n-1}\big)\right].$ (17)
Finally, the gradient of the loss function with respect to the $i$-th output feature map $\mathbf{y}_{i}^{(l-1)}$ of the $(l-1)^{\mathrm{th}}$ layer can be expressed as:
$\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{i}^{(l-1)}}$ $=\sum_{k=1}^{N_{l}}\boldsymbol{\delta}_{k}^{(l)}\odot\Big([\mathbf{S}_{k}^{(l)}]^{-1}\odot\frac{\partial\mathbf{R}_{k}^{(l)}}{\partial\mathbf{y}_{i}^{(l-1)}}$
$\quad-(\mathbf{R}_{k}^{(l)}\oslash[\mathbf{S}_{k}^{(l)}]^{\circ 2})\odot\frac{\partial\mathbf{S}_{k}^{(l)}}{\partial\mathbf{y}_{i}^{(l-1)}}\Big).$ (18)